calculate the H3O+ concentration for each pH

pH= 13

pH=2

pH=8
Answer #1

by definition

pH = -log [H30+]

[H30+] = 10-pH

1) given

pH = 13

so

[H30+] = 10-13

2) given

pH = 2

so

[H30+] = 10-2

3) given

pH = 8

so

[H30+] = 10-8